# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ Abstract
Model transfer presents a solution to the challenges of segmentation in the biomedical community, where the immense cost of data annotation is a major bottleneck in the use of deep learning. At the same time, hundreds of models get trained on biomedical data, submitted to challenges, and posted in model zoos and repositories. A major hurdle to wider adoption of pre-trained models lies in the lack of methods for best model selection. While such methods have been proposed for classification models, semantic and instance segmentation model ranking remain largely unaddressed, especially in a practically important setting where no labels are available on the target dataset. Similarly, if unsupervised domain adaptation is used, practitioners are faced with the task of selecting the best adapted model without target domain labels. Building on previous work linking model generalisation and consistency under perturbation, we propose the first unsupervised and source-free transferability estimator for semantic and instance segmentation tasks. We evaluate on multiple segmentation problems across biomedical imaging, finding a strong correlation between the rankings based on our estimator and rankings based on target dataset performance.
@@ 1 Introduction
Segmentation is a ubiquitous problem in biomedical image analysis as it plays a key role in the interpretation of biological and anatomical structures. While modern deep learning methods have significantly improved segmentation accuracy, their practical use remains limited by the need for labour-intensive, pixel-wise annotation of training data. Transfer learning [75, 45] aims to address this problem through re-use of networks trained on other datasets (“source”), either similar to the dataset of current interest (“target”) or sufficiently large to sample all possible targets. The latter approach has been prevalent in the analysis of natural images, while in biomedical imaging, many practitioners rely on training domain-specific models on their own data [28]. Community efforts such as the BioImage Model Zoo (BMZ) [49] make many such models publicly available, public challenges like the ToothFairy series [9] drive the creation and sharing of new models, and emerging large generalist models have had their first successes in the biomedical domain [3, 51, 64]. These initiatives raise challenges in selecting the best model for a target dataset when multiple models—trained on the same task but with varying settings, architectures, or source data—are available.
For the task of image classification, multiple supervised model selection/transferability ranking methods have been introduced [57, 16, 27, 80, 47, 74, 7, 36]. It has also been suggested that transferability ranking for semantic segmentation can be achieved as a simple extension of classification approaches [1]. In our experiments, this suggestion does not hold for biomedical imaging data, where classification-derived transferability metrics fail to reliably correlate with model performance on the target datasets. Furthermore, while semantic segmentation can be expressed as a pixel classification problem and thus derived from classification approaches, instance segmentation cannot. Instance segmentation outputs an unordered, variable‑length set of instances (masks, boxes) that only become discrete after model‑ and hyperparameter‑dependent instantiation steps (e.g., thresholding, binarisation, NMS/post‑processing). Instantiation results in a permutation invariant set of masks that have no stable, fixed‑dimensional correspondence between feature vectors, predictions, and target labels for the metrics to use. To the best of our knowledge, the problem of instance segmentation transferability ranking thus remains unaddressed, both for natural and for biomedical images. Supervised ranking exploiting target domain labels has other downsides: annotation, especially in 3D, can take weeks of expert time without any guarantee that a representative sample has been annotated. Unsupervised ranking is, therefore, much more practically useful. Also, recent advances in Unsupervised Domain Adaptation (UDA) have enabled fine-tuning without target domain labels, bringing forward the related problem of unsupervised evaluation of adapted model performance. First such methods have been introduced [59, 44, 78], again largely for classification models. However, in our experiments, the state-of-the-art for unsupervised validation [78] consistently underperforms for semantic segmentation. Also, like all methods directly extended from classification tasks, they are not applicable to instance segmentation.
Paper Contribution. The aim of this paper is to address the challenge of unsupervised segmentation model ranking to facilitate wider adoption of deep learning-based segmentation in the biomedical domain, where the difficulty of the annotation process and resource limitations hinder supervised training or model selection. Additionally, to ensure practical applicability within model repositories, we aim to operate in a source-free setting.
Building on the approaches proposed for model generalisation and uncertainty estimation, we (i) introduce a novel zero-shot transferability estimator for semantic and instance segmentation problems (Fig. 1). The estimator is based on prediction consistency under feature or input space perturbation and is both source-free and unsupervised. We (ii) evaluate the estimator on semantic and instance segmentation problems that represent highly popular tasks in microscopy: cell and nucleus segmentation in light microscopy and mitochondria segmentation in electron microscopy. We compare a diverse range of architectures and sizes, from small domain-specific models to generalist segmentation tools. Additionally, we (iii)
apply the estimator to models submitted to the ToothFairy challenge series [9], a benchmark for multi‑class 3D CBCT (Cone Beam Computed Tomography) semantic segmentation, reproducing challenge average DICE ranking without access to ground-truth. (iv) Beyond our main task of zero-shot transferability ranking, we apply our approach to rank models after UDA. (v) An extensive set of experiments shows that model rankings produced by our estimator are strongly correlated with target dataset performance, providing a simple but powerful model selection method enabling effective reuse of pre-trained models for practitioners in the field. Code will be made publicly available on acceptance.
@@ 2 Related Work
Qualitative approaches comparing datasets are common practice, but unreliable, as imperceptible differences in source and target dataset distributions, model architectures, and training regimes impact likely transfer success [62, 42, 45]. In the following, we review the related work in supervised transferability prediction and post-UDA model selection, highlighting work from the domain generalisation and uncertainty estimation fields that provided inspiration.
Supervised Transferability Metrics. Transferability metrics fall into two main categories: (1) label-comparison–based methods, such as LEEP [47], and (2) feature-embedding–based methods, which analyze the embedding space on target data with guidance from target labels. Examples include: LEEP [36], LogME [80], H-score [7], Regularised H-score [27], GBC [52], NCTI [74], and CC-FV [79], which, unlike the others, is explicitly designed for semantic segmentation. The above approaches require target domain labels and were mostly designed for classification. Although our proposed method is unsupervised, in the absence of unsupervised baselines, we select all these models as baselines for our experiments.
@@
Unsupervised Validation. Many popular UDA methods rely on self or adversarial training [32, 12] and are notoriously unstable, occasionally even leading to model degeneration [71, 73]. The domain of UDA validation, therefore, addresses a problem close to ours: how to select the best adapted model without target domain labels. The state-of-the-art in this field is the Transfer Score (TS) [78], developed for classification, but also applied to segmentation. It measures the classifier bias of a UDA model in the target domain along with the feature space transferability and discriminability. We use this score as a baseline in our experiments.
Domain Generalisation and Uncertainty Estimation. The task of model generalisation evaluation aims to train models that generalise better. While our goal is to estimate, rather than improve transferability, we can exploit similar approaches to measuring the generalisability, assuming that more generalisable models also transfer better. Here, consistency-based approaches [15, 61, 2] potentially satisfy our requirements for an unsupervised estimator and, although introduced for classification, could be extended to serve as a basis for semantic segmentation model ranking.
Early works [2, 61] suggest that prediction invariance under test-time-augmentation can be linked to model generalisability. However, they only focus on the in-domain (source) setting. In the case of [2] they rely on a carefully weighted penalty score to evaluate prediction invariance and [61] uses ground-truth to directly measure changes in the classification accuracy. Effective Invariance (EI) [15] was the first to address out-of-distribution (OOD) generalisability. EI measures both the consistency and confidence of probabilistic model outputs, showing that a combination of both is necessary for classification. However, it makes the method susceptible to failure for poorly calibrated models, which is especially relevant in the OOD transfer setting [50]. Furthermore, the EI cannot be extended to instance segmentation.
Consistency to perturbations has also been widely investigated in the uncertainty estimation field, where variance of prediction can be used as a measure of epistemic uncertainty [19, 39, 33, 43]. Uncertainty estimation aims to find well calibrated estimates of model confidence that correlate with true performance. The field is not concerned with transferability or generalisability of models and often makes use of supervised target labels to help calibrate uncertainty. Most relevant to our task are the post-hoc uncertainty estimators such as [39, 33, 43, 70] as ranking cannot influence model training. Rich theory links these approaches to Bayesian Neural Networks and Monte-Carlo Dropout [19], providing motivation for probing model responses under both input and feature space perturbations. As our ideal transferability estimator also needs to strongly correlate with model performance on target data, the work in these two domains provides inspiration for our approach.
@@ 3.1 Problem Definition
For unsupervised model transfer, a pre-trained model , trained on a supervised source domain is applied to an unlabelled target dataset taken from a target domain . Inference on can either be performed directly or after UDA, which further trains on an unlabelled subset of . We assume a transductive transfer setting [20], where the marginal distributions of the domains differ, , resulting in a domain shift, , but the source and target task are fixed (e.g., nuclei segmentation). We introduce an unsupervised consistency-based transferability metric (CTE) to correlate with the true performance (e.g., F1 score), which is not available without GT. Given a set of candidate source models , the goal of a transferability metric is to preserve the ranking if and only if .
3.2 Motivation
We propose to estimate model transferability based on the consistency of model outputs under perturbation. Focusing the analysis solely on model outputs enables a fully unsupervised and source-free transfer ranking. Since the output spaces of candidate models are fixed and directly comparable—unlike their feature spaces, which may differ substantially in composition and dimensionality across architectures—our approach provides a uniform basis for evaluation. Moreover, by operating in the final segmentation output space, our method allows, for the first time, direct assessment of the transferability of segmentation models, even across varied prediction processing/instantiation approaches.
The motivation for using consistency as a measure of transferability arises from the following considerations: a source model trained on a specific data distribution is expected to learn decision boundaries that partition this space by class. When applied to a target domain with a different distribution, these boundaries may fail to separate target classes effectively or may intersect clusters of target pixel embeddings. To assess transferability, we introduce small perturbations—either to the model input or its features—and observe whether pixel embeddings cross decision boundaries in the output prediction. Measuring such pixel class flips provides an assumption-free estimate of transferability, in contrast to prior methods that rely on the assumptions of feature-space compactness or separability, often modelled with Gaussian distributions [52, 36]. This is advantageous for semantic/instance segmentation, where common losses (e.g., combinations of Dice and cross-entropy) do not explicitly enforce feature compactness.
Prediction consistency can be viewed as a proxy for the margin of a model’s decision boundaries with respect to the target data. Models that transfer well are expected to exhibit larger margins, leading to better-separated classes and predictions that are less sensitive to perturbations. This intuition aligns with prior discussions on manifold smoothness and margin distribution [29, 46]. Related reasoning is also adopted in GBC [52], which uses class feature separability as a transferability metric. This metric, however, requires access to target labels, so we choose to probe for perturbation-induced class changes instead.
From this perspective, using input image perturbations, as customary in domain generalisation, or feature space perturbations, popular in uncertainty estimation, are conceptually equivalent. Unlike input perturbations for data augmentation in training, consistency perturbations need not represent realistic image changes between source and target distributions. Their purpose is to perturb pixel embeddings, testing the suitability of a model’s decision boundary for target data by measuring the number of output class label flips. Also, feature space perturbations do not necessarily correspond to realistic input image changes. In our experiments (Sec. 4), we evaluate several perturbation strategies and find all are effective.
3.3 Consistency Transferability Estimator (CTE)

### v3

# Unsupervised Source-Free Ranking of Biomedical Segmentation Models Under Distribution Shift

@@ Abstract
Model reuse offers a solution to the challenges of segmentation in biomedical imaging, where high data annotation costs remain a major bottleneck for deep learning. However, although many pretrained models are released through challenges, model zoos, and repositories, selecting the most suitable model for a new dataset remains difficult due to the lack of reliable model ranking methods. We introduce the first black-box-compatible framework for unsupervised and source-free ranking of semantic and instance segmentation models based on the consistency of predictions under perturbations. While ranking methods have been studied for classification and a few segmentation-related approaches exist, most target related tasks such as transferability estimation or model validation and typically rely on labelled data, feature-space access, or specific training assumptions. In contrast, our method directly addresses the repository setting and applies to both semantic and instance segmentation, for zero-shot reuse or after unsupervised domain adaptation. We evaluate the approach across a wide range of biomedical segmentation tasks in both 2D and 3D imaging, showing that our estimated rankings strongly correlate with true target-domain model performance rankings.
@@ 1 Introduction
Segmentation is a ubiquitous problem in biomedical images, vital for the interpretation of biological and anatomical structures. While modern deep learning has significantly improved segmentation accuracy, practical use is limited by the labour-intensive annotation of training data. Model transfer [weiss_survey_2016, mustafa_supervised_2021], with or without fine-tuning, offers a potential solution by reusing pre-trained “source” networks for new “target” datasets of interest. In biomedical imaging, where comprehensive public training datasets are rare, and new imaging modalities continue to emerge, practitioners often train domain-specific models [isensee2021nnu]. Community initiatives such as the BioImage Model Zoo (BMZ) [ouyang_bioimage_2022], public challenges such as ToothFairy [2024TMI], and emerging large generalist models [archit_segment_2025, pachitariu_cellpose-sam_2025, stringer_cellpose_2020, wasserthal_totalsegmentator_2023] have created a growing ecosystem of reusable models. This raises the practical challenge of model selection when multiple candidate models—trained on the same task but with varying settings, architectures, or source data—are available.
Transferability estimation [agostinelli_how_2022] aims to address this task, often in a source-free, model-agnostic setting appropriate for repository model ranking. Most prior work addresses the task of image classification, where multiple supervised methods have been introduced [renggli_which_2020, ding_which_2024, ibrahim_newer_2023, you_logme_2021, nguyen_leep_2020, wang_how_2023, bao_information-theoretic_2022, li_ranking_2020] to rank post-fine-tuning performance. It has also been suggested that transferability ranking for semantic segmentation can be achieved by simple extension of classification approaches [agostinelli_how_2022]. In our experiments, this claim does not hold for biomedical imaging data, where classification-derived methods fail to reliably correlate with model target data performance. Furthermore, while semantic segmentation can be expressed as pixel classification, instance segmentation cannot. Instance segmentation outputs a permutation invariant set of labels that arise from model‑dependent post-processing (e.g., connected components analysis), so the labels have no stable, fixed‑dimensional correspondence to instance representation in feature space. The problem of instance segmentation transferability ranking remains largely unaddressed for both natural and biomedical images.
While supervised transferability estimation can save fine-tuning costs, annotation, especially in 3D, can take weeks of expert time without guaranteeing that a representative set has been labelled. Unsupervised ranking, therefore, has significant practical advantages. Also, recent advances in Unsupervised Domain Adaptation (UDA) enable potential improvements to model target data performance without labels, but still require unsupervised ranking of the adapted models. First such methods have been introduced [saito_tune_2021, morerio_minimal-entropy_2017, yang_can_2024], again largely for classification. However, in our experiments, the state-of-the-art (SOTA) for UDA validation [yang_can_2024] fails to reliably rank semantic segmentation models and, like other classification-derived approaches, cannot be applied to instance segmentation.
While model repositories exist—HuggingFace alone hosts over 2.7 million models—practical ranking methods remain scarce because real-world deployment requires them to be unsupervised, source-free, and model-agnostic—constraints that few existing approaches satisfy. To address this gap, we therefore consider the problem of selecting the best segmentation model from a repository for a new dataset without access to target labels.
Contribution. In summary, our contribution is fourfold:
We introduce the first unsupervised, source-free method for ranking both semantic and instance segmentation models based on prediction consistency under feature- or input-space perturbations (Fig.˜1);
The method is training- and architecture-agnostic and operates only on model predictions. It can be applied to black-box models, making it compatible with heterogeneous model zoos even if using inference APIs or containerised deployments;
We evaluate on realistic biomedical transfer scenarios spanning nuclei, cell, and mitochondria segmentation in light and electron microscopy. Our method successfully ranks models ranging from domain-specific architectures to large generalist segmentation models (e.g., SAM [archit_segment_2025] and CellposeSAM [pachitariu_cellpose-sam_2025]), both for direct application and after unsupervised domain adaptation;
We further demonstrate the practical applicability of our approach by ranking submissions from the ToothFairy challenge [2025CVPR], a benchmark for multi-class 3D CBCT semantic segmentation. Without access to test dataset labels, our method successfully reproduces the challenge’s average model DICE ranking.
@@ 2 Related Work
In the following, we review supervised transferability metrics, UDA validation, performance estimation, generalisation, and uncertainty estimation while highlighting the distinct challenges of unsupervised model ranking. These fields address distinct yet related problems. While they share many methodological approaches, they use different terminology and rarely cross-cite.
Supervised Transferability Metrics. Transferability metrics (TM) aim to rank a set of models by their post-fine-tuned target performance. Existing work follows two main categories: (i) label-comparison-based methods, such as LEEP [nguyen_leep_2020], and (ii) feature-embedding-based methods, which analyse the target data embedding space with guidance from target labels (LEEP [li_ranking_2020], LogME [you_logme_2021], H-score [bao_information-theoretic_2022], Regularised H-score [ibrahim_newer_2023], GBC [pandy_transferability_2021], NCTI [wang_how_2023]). They all require target domain labels and were initially designed for classification, except CC-FV [yang_pick_2023], which focuses on semantic segmentation. Our task and method are unsupervised, but fundamentally aim to answer a very similar question of selecting the most suitable source models for a given domain-shifted target dataset [agostinelli_how_2022]. Hence, many TM methods satisfy our requirements of being source-free and model agnostic and can serve as strong baselines for ranking ‘zero-shot transferability’. Our key difference: label-free, no fine-tuning assumption.
@@
UDA Validation. Many popular UDA methods rely on self or adversarial training [kumar_understanding_2020, cao_multi-modal_2023] and are notoriously unstable [wang_deep_2018, wang_characterizing_2019]. Hence, UDA validation addresses a problem close to ours: how to select the best adapted model without target domain labels. The SOTA in this field is the Transfer Score (TS) [yang_can_2024], developed for classification, but also tested on segmentation. It measures the post-UDA model classifier bias in the target domain along with the feature space transferability and discriminability. We use TS as a further baseline in our experiments. Our key difference: ranking across heterogeneous models, not only UDA setting.
Performance Estimation. Another related field is unlabelled performance estimation (PE), also known as quality control (QC), where the aim is to directly estimate performance metrics. However, in PE, the unit of analysis is typically a single model, and methods aim to assess the quality of the model’s predictions in deployment, e.g., under distribution drift. In contrast, our method estimates the relative success of a candidate set of models applied to specific new domain-shifted target data. Thus we do not require the score to equate to a direct performance metric for each model, as we aim to recover a relative ranking across the model set. PE methods need to rely on assumptions about the nature of the domain shift to uniquely identify the target conditional [garg_leveraging_2022]. At the same time, PE methods are not required to be target label-free [valindria_reverse_2017], source-free [garg_leveraging_2022, bialek_estimating_2025] or model-training agnostic [lin_novel_2023, hoebel_exploration_2020, roy_bayesian_2019, robinson_real-time_2018, kivimaki_confidence-based_2025, mehrtash_confidence_2020], making most of them inapplicable for model selection in the unsupervised model repository setting. Using a recent benchmark [yu_odp-bench_2025], we identify a few exceptions that satisfy our requirements, namely Nuclear Norm [deng_confidence_2023] and Dispersion [xie_importance_2023]. The first seeks to quantify the confidence and dispersity of model predictions via the nuclear norm, while the second, similarly to [pandy_transferability_2021, yang_pick_2023] seeks to use inter-class feature dispersion as a surrogate for model performance. We implement both as baselines. Our key difference: repository ranking vs single-model QC, source-free, model-training agnostic.
Ensemble-based PE methods [sims_seg_2023, baek_agreement---line_2023] incur a large computational overhead, but can in principle be applied to ranking, with [sims_seg_2023] specifically proposed for segmentation. However, [sims_seg_2023] reduces the evaluation to object centroids, making the method insensitive to object shape, especially for irregularly shaped objects. Missing pixelwise object evaluation, the method is closer to object detection than instance segmentation ranking, but still provides a useful baseline as the only other approach to address instance segmentation.
Generalisation and Uncertainty Estimation. Generalisability is another term used to describe the ability of a model to perform across new datasets. While often used for in-domain (ID) data, out-of-distribution (OOD) generalisability also exists. Typically, it is considered as an intrinsic model property that can be improved, e.g., by learning more robust and broadly applicable features. Our aim is not to improve target performance, yet we can exploit similar approaches to those measuring generalisability, but in a specific paired model-dataset setting. Ranking differs from generalisability as it explicitly measures the compatibility between a model and a target dataset either directly or post-UDA and is not a property of the model alone. Consistency-based approaches [deng_strong_2022, schiff_predicting_2021, k_robustness_2021] provide potentially suitable unsupervised methods and have been explored for classification. Early works [k_robustness_2021, schiff_predicting_2021] suggest that prediction invariance under test-time-augmentation (TTA) can be linked to model generalisability, but do not explore feature perturbations. Furthermore, they only focus on the ID setting, relying on carefully weighted penalty scores [k_robustness_2021], or directly using ground-truth [schiff_predicting_2021] to evaluate prediction invariance. Effective Invariance (EI) [deng_strong_2022] addresses OOD generalisability, measuring both the consistency and confidence of probabilistic model outputs. However, EI cannot be extended to instance segmentation.
Consistency to perturbations has also been investigated in the uncertainty estimation (UE) field, where rich theory ties the variance of predictions to an approximation of Bayesian neural networks (BNNs) [gal_dropout_2016, loquercio_general_2020, ledda_dropout_2023, mi_training-free_2022, wang_aleatoric_2019, wang_epistemic_2024]. Calibrating model confidence to correlate with true performance is a key challenge and often requires access to source/target labelled validation sets [loquercio_general_2020, ledda_dropout_2023] or restriction of the training procedure [gal_dropout_2016]. The UE field itself is not concerned with assessing model applicability to target data, but rather aims to improve model interpretability via the quantification of uncertainty. However, UE has been widely used in many PE, QC, or failure detection methods [zenk_comparative_2025, lin_novel_2023, devries_leveraging_2018, hoebel_exploration_2020, kivimaki_confidence-based_2025, mehrtash_confidence_2020, jungo_analyzing_2020, roy_bayesian_2019] that explicitly conflate estimating performance and uncertainty, thus suffering the same limitations. These approaches are typically developed to assist medical experts in segmentation and provide per-image evaluation of a single model’s predictions. While important, such methods are not suitable for model selection in model zoos, adding unnecessary, hard to satisfy uncertainty calibration requirements (i.e., restricting the model training [devries_leveraging_2018, hoebel_exploration_2020, roy_bayesian_2019, kivimaki_confidence-based_2025, mehrtash_confidence_2020]). Our key difference: no uncertainty calibration, semantic and instance segmentation, model-agnostic.
Following source-free, model-agnostic, unsupervised constraints of a model repository setting, we use the building blocks from the above fields to create a ranking-specific approach, stripping away complexities associated with confidence calibration or feature space reasoning, and explicitly do not use perturbations for UE, as described in Sec.˜3. Our method, for the first time, can be applied in a unified framework for both semantic and instance segmentation ranking.
@@ 3 Methods
We rank model performance via the consistency of model outputs under perturbation. Focusing on output analysis enables a fully unsupervised and source-free ranking. Since the output spaces of candidate models are fixed and directly comparable—unlike their feature spaces, which may differ in composition and dimensionality—our approach provides a uniform evaluation basis. Moreover, our method, for the first time, directly assesses pixel-wise instance segmentation performance across arbitrary instantiation methods. SEG [sims_seg_2023] has previously attempted to address instance segmentation, but reduces the problem to object detection based on centroids. This constrains SEG to circular objects and disregards pixel-wise information. We motivate our consistency-based approach as follows: a source model learns decision boundaries to partition data by class. When applied to a distribution-shifted target domain, these boundaries may fail to separate classes effectively or may intersect embedding clusters. To assess model-dataset compatibility, we introduce small perturbations—either to the model input or its features—and observe predicted class flips as pixel embeddings cross decision boundaries. Measuring such pixel class flips provides a model-agnostic estimate of compatibility, free from feature space-based reasoning used in prior methods (e.g., feature-space compactness or separability [pandy_transferability_2021, li_ranking_2020, xie_importance_2023, deng_confidence_2023, wang_how_2023, yang_can_2024, yang_pick_2023]). This is advantageous for semantic/instance segmentation, where common losses (e.g., Dice and cross-entropy) do not explicitly enforce feature compactness.
Prediction consistency can be viewed as a proxy for the margin of a model’s decision boundaries on target data. Models that transfer well are expected to exhibit larger margins, leading to better-separated classes and less perturbation sensitive predictions. This intuition aligns with prior discussions on manifold smoothness and margin distribution [jiang_predicting_2018, ng_predicting_2023]. Without access to labels we rely on perturbation-induced class changes to indicate separability. From this perspective, input-image perturbations (commonly in generalization) and feature-space perturbations (popular in uncertainty estimation) are conceptually equivalent.
Consistency perturbations need not represent realistic image changes between source and target distributions; e.g., feature-space perturbations do not necessarily have realistic image interpretations. Their purpose is to perturb pixel embeddings, testing the suitability of a model’s decision boundary for target data by measuring the number of output class label flips. We explicitly do not use perturbations to estimate uncertainty, and do not use MC-dropout [gal_dropout_2016], which is used in many PE methods for calibrated confidence scores and must be applied during both training and inference. Instead, we employ test-time-dropout (TTD) [tompson_efficient_2015, ledda_dropout_2023] which can be applied to arbitrary layers at inference time only. Thus remaining model agnostic and setting no model training requirements.
@@ 3.1 Problem Definition
Considering a pre-trained model , trained on a source domain we apply to an unlabelled target dataset , sampled from a shifted target domain . Inference on can be performed directly or after UDA, which further trains on a subset of . We enforce that the source and target task are the same (e.g., nuclei segmentation). Given a set of candidate source models , our goal is to rank the models on so that the ranking correlates with true performance (e.g., F1 score) ranking, which is not available without labels. We introduce an unsupervised consistency-based model ranking method (CMR) that aims to preserve the ranking if and only if .
3.2 Our Consistency Model Ranking (CMR)